### 29 Mar 2026 v1

@@ 4.5 Ablation Study
Figure 3: ACD convergence analysis across 111 MM-Bench problems.
4.6 Convergence Behavior of ACD
@@ 5 Discussion and Conclusion
We presented Sci-Mind, a cognitively-inspired framework that integrates Experiential Memory Recall, Adversarial Cognitive Dialectic, and Self-Validating Execution Strategy for autonomous mathematical modeling. Experiments on MM-Bench and EngiBench across four backbones confirm state-of-the-art performance in both modeling rigorousness and code executability, with gains orthogonal to backbone capabilities. Our analysis reveals two broader insights: objective asymmetry matters more than agent count for effective verification, as the gap between Homogeneous Debate and full ACD far exceeds the gap between Self-Reflection and Homogeneous Debate; and the abstraction-implementation gap is a distinct failure mode where agents possess declarative knowledge to formulate correct models but lack procedural knowledge to implement them, which retrieval of executable structural priors can directly address.
One limitation is that our experiments focus on competition-style modeling tasks, and generalizability to industrial or laboratory scientific workflows with proprietary data and domain-specific constraints remains to be validated. Looking forward, extending the fixed verification predicates to a learnable generation mechanism, and accumulating domain-specific critique templates through finer-grained self-evolution represent promising directions. We hope Sci-Mind provides a useful step toward autonomous agents that reason about scientific problems with the rigor and groundedness of human domain experts.

### 02 Apr 2026 v2

@@ 4.3 Cross-Domain Generalization
4.4 Convergence Behavior of ACD
Figure 3: ACD convergence analysis across 111 MM-Bench problems.
4.5 Ablation Study
@@ 5 Discussion and Conclusion
We presented Sci-Mind, a cognitively-inspired framework that integrates Experiential Memory Recall, Adversarial Cognitive Dialectic, and Self-Validating Execution Strategy for autonomous mathematical modeling. Experiments on MM-Bench and EngiBench across four backbones confirm state-of-the-art performance in both modeling rigorousness and code executability, with gains orthogonal to backbone capabilities.
Our analysis reveals two broader insights. First, objective asymmetry matters more than agent count for effective verification: the gap between Homogeneous Debate and full ACD far exceeds that between Self-Reflection and Homogeneous Debate, suggesting that decomposing verification into competing objectives is more effective than scaling homogeneous reviewers. Second, the abstraction-implementation gap is a distinct failure mode where agents can formulate correct models but cannot implement them, which retrieval of executable structural priors directly addresses.
One limitation is that our experiments focus on competition-style modeling tasks, and generalizability to industrial or laboratory workflows remains to be validated. This motivates two future directions: extending the fixed verification predicates to a learnable mechanism that adapts to new domains, and accumulating domain-specific critique templates through finer-grained self-evolution to accelerate convergence on specialized tasks. We hope Sci-Mind provides a useful step toward autonomous agents that reason about scientific problems with the rigor and groundedness of human domain experts.